the value of [1+cospi/6] [1+cospi/3] [1+cos2pi/3] [1+cos7pi/6] is
a).3/16 b). 3/8 c).3/4 d). 1/2

Grade:11

1 Answers

Sankalp Das
6 years ago
=[1+cospi/6] [1+cospi/3] [1+cos2pi/3] [1+cos7pi/6] =(1+√3/2)(1+1/2)(1-1/2)(1-√3/2)=(1-3/4)(1-1/4)=1/4×3/4=3/16C is correct